# Identifying the impact of local connectivity patterns on dynamics in excitatory-inhibitory networks

### v2

@@ IV.2.2 Paradoxical responses
IV.3 Sparse excitatory-inhibitory networks
Figure 8: Impact of chain motifs on population-averaged mean of entries on connectivity vectors for sparse networks (red and blue) and their Gaussian approximations (orange and green). (a, b) Population-averaged mean values of entries $m_{i}^{(r)p}$ on the right connectivity vectors. Subplots (a) and (b) respectively show the mean values for $p=E$ and $p=I$. (c, d) Population-averaged mean values of entries $n_{i}^{(r)p}$ on the left connectivity vectors. Subplots (c) and (d) respectively show the mean values for $p=E$ and $p=I$. The insets show the distribution of the mean values of the elements in the connectivity vectors corresponding to the eigenvalues in the bulk (red: excitatory population; blue: inhibitory population). An additional eigenvalue emerges from the bulk $\tau^{c}_{EE}=0.11$. For other network parameters see TABLE 2.
Figure 9: Impact of chain motifs on the mean response $\chi_{E}$ of excitatory neurons to uniform external inputs in sparse networks, as a function of strength of chain motifs, which is heterogeneous across blocks of the connectivity matrix, with $\tau^{c}=\tau_{EE}^{c}=\tau_{II}^{c}=-\tau_{EI}^{c}=-\tau_{IE}^{c}$. The asterisks with error bars show the mean and s.d. of $\chi_{E}$ in 30 realizations of the full sparse network. The triangles with error bars show the mean and s.d. of $\chi_{E}$ in 30 realizations of the equivalent Gaussian approximation. The lines represent the predictions of the low-rank approximation in Eq. (11), where the $a_{x}^{p}$ were obtained numerically by diagonalizing the sparse connectivity matrix. The dashed line is the prediction obtained from the mean connectivity $\mathbf{J}^{0}$ (Eq. (51)). The colored light lines illustrate the respective contributions of the two unit-rank terms in the low-rank approximation of the response of the sparse networks. Orange and green denote respectively the modes corresponding to the negative and positive outliers. Inset: outlying eigenvalues as a function of $\tau^{c}$, compared to the bulk of the radius. The positive outlier emerges from the bulk for $\tau^{c}=0.11$ (vertical dash-dot line). For other network parameters, see TABLE 2.
@@ IV.3 Sparse excitatory-inhibitory networks
Similarly to fully-connected networks, increasing the overall strength of chain motifs by increasing $\rho^{c}$ amplifies the response $\chi_{p}$ to uniform inputs across the network (Fig 9). This effect is well captured by a rank-two approximation, which shows that the amplification is also here due to the positive outlier emerging from the bulk of the eigenspectrum (Fig 9 colored lines).
@@
Inspecting the mean responses $\chi_{pq}$ of population $p$ to uniform inputs in population $q$ however revealed important differences between sparse networks (Figs. 10,D.4) and fully connected networks (Fig 7). Indeed, in sparse networks, the mean response of individual populations appeared to change in a direction opposite to fully connected networks when the strength of chain motifs is increased by increasing $\rho^{c}$. In particular, the mean response of inhibitory neurons to its inputs increases with $\tau^{c}:=\tau_{EE}^{c}=\tau_{II}^{c}=-\tau_{EI}^{c}=-\tau_{IE}^{c}$ (Fig 10(d)), while in fully-connected networks it decreases (Fig 7(d)). In sparse networks chain motifs can therefore induce a transition from paradoxical inhibitory responses to non-paradoxical responses ($\chi_{II}$ in Fig. D.4), while in fully-connected networks the situation is the opposite. This difference can be traced back to the mean values $a_{m_{r}}^{p}$ and $a_{n_{r}}^{p}$ of the entries of connectivity vectors that determine the responses $\chi_{pq}$ through Eq. 48. Specifically, the mean values $a_{n_{2}}^{E}$ and $a_{n_{2}}^{E}$ of the second connectivity vector onto the excitatory and inhibitory populations have opposite signs in sparse networks (Figs D.38 (c,d)) compared to fully-connected networks (Fig 5 (c,d)).
@@
Figure 11: Impact of the heterogeneity of chain motifs on the response to external inputs. (a) Response $\chi_{II}$ of inhibitory neurons as a function of $\tau^{c}_{EE/II}=\tau^{c}_{EE}=\tau^{c}_{II}$ and $\tau^{c}_{EI/IE}=\tau^{c}_{EI}=\tau^{c}_{IE}$, for the fully connected Gaussian approximation of sparse networks. The green line corresponds to homogeneous chain motifs ($\tau^{c}_{EE}=\tau^{c}_{II}=\tau^{c}_{EI}=\tau^{c}_{IE}$), for which the response is shown in Fig 7 . The purple line corresponds to homogeneous chain motifs occurence $\rho^{C}$, leading to $\tau^{c}_{EE}=\tau^{c}_{II}=-\tau^{c}_{EI}=-\tau^{c}_{IE}$, for which the response is shown in Fig 10. The yellow line shows the prediction for networks with $\tau_{EI/IE}^{c}=0$. (b) Response function $\chi_{EI}$ and (c) response function $\chi_{II}$ in sparse networks and corresponding fully connected Gaussian approximations with $\tau_{EI/IE}^{c}=0$. Details of subplots (b) and (c) are identical to Fig 10. The other network parameters are identical to Fig 10.
IV.4 Effects of chain motif heterogeneity on paradoxical responses
@@
To understand the origin of seemingly opposite effects of chain motifs on paradoxical responses in sparse and fully-connected networks, we note that in fully-connected networks we used homogeneous, positive $\tau^{c}$, while for sparse networks with a uniform chain motif occurrence $\rho^{c}$, the correlation coefficients are different across blocks of the connectivity matrix, with $\tau_{EE}^{c}=\tau_{II}^{c}$ positive and $\tau_{EI}^{c}=\tau_{IE}^{c}$ negative. We hypothesized that this difference may lead to the distinct dependences of the responses $\chi_{pq}$ to the strength of chain motifs. To investigate this possibility, we computed the responses $\chi_{II}$ in Gaussian networks where the chain motif strengths were not homogeneous across blocks of the connectivity matrix, but were instead determined by two parameters $\tau_{EE/II}^{c}$ and $\tau_{EI/EI}^{c}$ such that $\tau_{EE}^{c}=\tau_{II}^{c}=\tau_{EE/II}^{c}$ and $\tau_{EI}^{c}=\tau_{IE}^{c}=\tau_{EI/EI}^{c}$ (Fig 11). This analysis reconciled the findings in Fig 7 and Fig 10: increasing jointly $\tau_{EE/II}^{c}=\tau_{EI/IE}^{c}$ leads to a decreasing response $\chi_{II}$(green line in Fig 11), while setting $\tau_{EE/II}^{c}=-\tau_{EI/IE}^{c}$ instead leads to an increase of $\chi_{II}$ with the strength of chain motifs (purple line in Fig 11). We moreover predicted that setting $\tau_{EI/IE}^{c}=0$ and increasing $\tau_{EE/II}^{c}$ would lead to a decreasing $\chi_{II}$, and therefore a switch from a non-paradoxical to a paradoxical response (yellow line and frame in Fig 11). We directly verified this prediction in sparse networks by setting $\rho^{c}_{EI/IE}=c^{2}$ and therefore zero correlation coefficients $\tau_{EI/IE}^{c}=0$, while maintaining $\tau_{EE/II}^{c}=\tau^{c}>0$ and gradually increasing this value (Fig 11(b) and (c)).
Altogether, our results show that the nature of the response to external inputs strongly, and via explicitly predictable trends, depends on the presence of chain motifs and their distribution across populations.
@@ V Discussion
Our findings demonstrate that correlations between weights of different synapses, and in particular chain motifs, can exert a dominant impact on the dynamics in networks of recurrently connected neurons. Specifically, our mathematical analyses reveal that even a weak over-representation of chain motifs leads to the emergence of isolated eigenmodes distinct from networks with uncorrelated synaptic weights. In inhibition-dominated networks, these additional eigenmodes are associated with a positive eigenvalue that reflects positive feedback induced by chain motifs. This positive feedback leads to a potential instability that requires revisiting the classical balance conditions.
@@
The positive eigenmode generated by chain motifs strongly contributes to population-averaged responses to external inputs. In particular, we show that through this eigenmode, chain motifs can induce paradoxical responses in networks where such responses are not expected based on mean connectivity alone [10, 15, 19, 60, 61]. Conversely, in inhibition-stabilized networks that exhibit paradoxical responses in the absence of correlations between synapses, chain motifs can shift the dynamics towards non-paradoxical responses. Mean connectivity therefore does not predict the nature of the responses when chain motifs are present, but we expect that paradoxical responses still correspond to networks where the excitatory sub-network is unstable by itself [15].
@@
Previous works have examined the effects of connectivity motifs on the continuous bulk of the spectrum [49, 32, 30, 31, 62]. Here we instead focus on eigenmodes corresponding to eigenvalues that lie outside of the bulk. Our results show that connectivity motifs have distinct effects on these outliers, which determine the responses to external inputs.
@@
For the sake of analytical tractability, our study is based on a number of simplifying assumptions. We considered only linear, rate-based dynamics, while a number of previous studies outlined the effects of the non-linearity on paradoxical responses, in particular in supralinear-stabilized networks of excitatory and inhibitory neurons [14, 13, 12, 63, 64]. Although a full non-linear analysis of N-dimensional networks with correlated synapses is in general untractable, networks with low-rank connectivity allow for an exact low-dimensional reduction even in the presence of non-linearities [42, 65, 66]. Applying the low-rank approximation of correlated connectivity to supralinear stabilized networks is therefore an interesting direction for future research. Beyond rate-based models, numerical analyses of integrate-and-fire networks with low-rank connectivity structure suggest that the analytical insights from rate networks extend to spiking networks [67].
@@
Altogether, our results argue for the importance of moving beyond the classical paradigm of excitatory-inhibitory networks with uncorrelated synaptic weights. These models have been instrumental for generating mechanistic interpretations of neural recordings, in particular for spontaneous activity [7, 8, 58, 57] and responses to external inputs [68, 11, 13, 14, 21, 20, 69]. Yet they strongly rely on the assumption that the strengths of different synapses are statistically independent and therefore determined by their mean and variance. Experimental recordings of synaptic weights clearly show that individual synapses are not independent, and our study demonstrates that basic correlations can have strong effects on dynamics, both in terms of stability and the nature of responses to inputs. More generally, the computations performed by networks of neurons in the brain stem from the structure in the connections at both the large scale of cell-type averages and the smaller scale of synaptic configurations beyond these averages, and the analytical approach developed here offers a way of combining these levels of description to advance the links between connectivity structure and the dynamics of computation.

### v3

@@ IV.3.2 Paradoxical responses
We next examine the paradoxical responses from the perspective of effective connectivity, which in this case has a unit-rank structure, and is given by
$J^{eff}_{ij}=\left\{\begin{matrix}J^{0}+\sigma^{2}\tau^{c},&j\in N_{E}\\ -gJ^{0}+\sigma^{2}\tau^{c},&j\in N_{I}.\ \end{matrix}\right.$ (55)
Using this effective connectivity and Eq. (10), we accurately predict the response function $\chi_{p},~{}\chi_{pq}$ of the full network (orange dashed lines in Figs. 6 and 7).
Importantly, for an equivalent deterministic network defined by $\mathbf{J}^{eff}$, the paradoxical response of the inhibitory neuron population is determined by the stability of the E-E subnetwork $\mathbf{J}^{eff}_{EE}$ [15], which is made up only of synapses from excitatory to excitatory neurons. The only non-zero eigenvalue of this subnetwork is $\lambda_{EE}^{eff}=N_{E}(J^{0}+\sigma^{2}\tau^{c})$. The inhibitory population will respond paradoxically if this eigenvalue is greater than 1 and the E-E subnetwork becomes unstable [15]. Therefore, for $N_{E}J^{0}<1$, the eigenvalue of the E-E subnetwork crosses 1 for $\tau^{c}=(\frac{1}{N_{E}}-J^{0})/\sigma^{2}$, and this criterion predicts well the appearance of paradoxical inhibitory responses in the full network (Fig. 7, black triangle).
IV.4 Sparse excitatory-inhibitory networks
Figure 8: Impact of chain motifs on population-averaged mean of entries on connectivity vectors for sparse networks (red and green) and their Gaussian approximations (orange and blue). (a, b) Population-averaged mean values of entries $m_{i}^{(r)p}$ on the right connectivity vectors. Subplots (a) and (b) respectively show the mean values for $p=E$ and $p=I$. (c, d) Population-averaged mean values of entries $n_{i}^{(r)p}$ on the left connectivity vectors. Subplots (c) and (d) respectively show the mean values for $p=E$ and $p=I$. The insets show the distribution of the mean values of the elements in the connectivity vectors corresponding to the eigenvalues in the bulk (red: excitatory population; blue: inhibitory population). An additional eigenvalue emerges from the bulk $\tau^{c}_{EE}=0.11$. For other network parameters see TABLE 2.
Figure 9: Impact of chain motifs on the mean response $\chi_{E}$ of excitatory neurons to uniform external inputs in sparse networks, as a function of strength of chain motifs, which is heterogeneous across blocks of the connectivity matrix, with $\tau^{c}=\tau_{EE}^{c}=\tau_{II}^{c}=-\tau_{EI}^{c}=-\tau_{IE}^{c}$. The asterisks with error bars show the mean and s.d. of $\chi_{E}$ in 30 realizations of the full sparse network. The triangles with error bars show the mean and s.d. of $\chi_{E}$ in 30 realizations of the equivalent Gaussian approximation. The lines represent the predictions of the low-rank approximation in Eq. (11), where the $a_{x}^{p}$ were obtained numerically by diagonalizing the sparse connectivity matrix. The gray dashed line is the prediction obtained from the mean connectivity $\mathbf{J}^{0}$ (Eq. (52)), the orange dashed line is the prediction obtained from the effective connectivity $\mathbf{J}^{eff}$. The colored light lines illustrate the respective contributions of the two unit-rank terms in the low-rank approximation of the response of the sparse networks. Red and green denote respectively the modes corresponding to the negative and positive outliers. Inset: outlying eigenvalues as a function of $\tau^{c}$, compared to the bulk of the radius. The positive outlier emerges from the bulk for $\tau^{c}=0.11$ (vertical dash-dot line). For other network parameters, see TABLE 2.
@@ IV.4 Sparse excitatory-inhibitory networks
Similarly to fully-connected networks, increasing the overall strength of chain motifs by increasing $\rho^{c}$ amplifies the response $\chi_{p}$ to uniform inputs across the network (Fig 9). This effect is well captured both by the effective connectivity and by a rank-two approximation, which shows that the amplification is also here due to the positive outlier emerging from the bulk of the eigenspectrum (Fig 9 colored lines).
@@
Inspecting the mean responses $\chi_{pq}$ of population $p$ to uniform inputs in population $q$ however revealed important differences between sparse networks (Figs. 10,E.4) and fully connected networks (Fig 7). Indeed, in sparse networks, the mean response of individual populations appeared to change in a direction opposite to fully connected networks when the strength of chain motifs is increased by increasing $\rho^{c}$. In particular, the mean response of inhibitory neurons to its inputs increases with $\tau^{c}:=\tau_{EE}^{c}=\tau_{II}^{c}=-\tau_{EI}^{c}=-\tau_{IE}^{c}$ (Fig 10(d)), while in fully-connected networks it decreases (Fig 7(d)). In sparse networks chain motifs can therefore induce a transition from paradoxical inhibitory responses to non-paradoxical responses ($\chi_{II}$ in Fig. E.4), while in fully-connected networks the situation is the opposite. This difference can be traced back to the mean values $a_{m_{r}}^{p}$ and $a_{n_{r}}^{p}$ of the entries of connectivity vectors that determine the responses $\chi_{pq}$ through Eq. 48. Specifically, the mean values $a_{n_{2}}^{E}$ and $a_{n_{2}}^{E}$ of the second connectivity vector onto the excitatory and inhibitory populations have opposite signs in sparse networks (Figs E.38 (c,d)) compared to fully-connected networks (Fig 5 (c,d)).
@@
Figure 11: Impact of the heterogeneity of chain motifs on the response to external inputs. (a) Response $\chi_{II}$ of inhibitory neurons as a function of $\tau^{c}_{EE/II}=\tau^{c}_{EE}=\tau^{c}_{II}$ and $\tau^{c}_{EI/IE}=\tau^{c}_{EI}=\tau^{c}_{IE}$, for the fully connected Gaussian approximation of sparse networks. The green line corresponds to homogeneous chain motifs ($\tau^{c}_{EE}=\tau^{c}_{II}=\tau^{c}_{EI}=\tau^{c}_{IE}$), for which the response is shown in Fig 7 . The brown line corresponds to homogeneous chain motifs occurence $\rho^{C}$, leading to $\tau^{c}_{EE}=\tau^{c}_{II}=-\tau^{c}_{EI}=-\tau^{c}_{IE}$, for which the response is shown in Fig 10. The yellow line shows the prediction for networks with $\tau_{EI/IE}^{c}=0$. Coordinates (squares) with a dashed border indicate the first value of $\tau_{EI/IE}^{c}$ that results in a paradoxical effect as $\tau_{EI/IE}^{c}$ gradually increases, while under a fixed $\tau_{EE/II}^{c}$. The orange dashed line represents the theoretical prediction for the boundary between the paradoxical and the non-paradoxical regime using $\mathbf{J}^{eff}$ and Eq. (10), the orange solid line represents the prediction using the stability of the E-E subnetwork $\mathbf{J}^{eff}_{EE}$ (Eq. (58)) (b) Response function $\chi_{EI}$ and (c) response function $\chi_{II}$ in sparse networks and corresponding fully connected Gaussian approximations with $\tau_{EI/IE}^{c}=0$. Details of subplots (b) and (c) are identical to Fig 10. The other network parameters are identical to Fig 10.
IV.5 Effects of chain motif heterogeneity on paradoxical responses
@@
To understand the origin of seemingly opposite effects of chain motifs on paradoxical responses in sparse and fully-connected networks, we note that in fully-connected networks we used homogeneous, positive $\tau^{c}$, while for sparse networks with a uniform chain motif occurrence $\rho^{c}$, the correlation coefficients are different across blocks of the connectivity matrix, with $\tau_{EE}^{c}=\tau_{II}^{c}$ positive and $\tau_{EI}^{c}=\tau_{IE}^{c}$ negative. We hypothesized that this difference may lead to the distinct dependences of the responses $\chi_{pq}$ to the strength of chain motifs. To investigate this possibility, we computed the responses $\chi_{II}$ in Gaussian networks where the chain motif strengths were not homogeneous across blocks of the connectivity matrix, but were instead determined by two parameters $\tau_{EE/II}^{c}$ and $\tau_{EI/EI}^{c}$ such that $\tau_{EE}^{c}=\tau_{II}^{c}=\tau_{EE/II}^{c}$ and $\tau_{EI}^{c}=\tau_{IE}^{c}=\tau_{EI/EI}^{c}$ (Fig 11). This analysis reconciled the findings in Fig 7 and Fig 10: increasing jointly $\tau_{EE/II}^{c}=\tau_{EI/IE}^{c}$ leads to a decreasing response $\chi_{II}$(green line in Fig 11), while setting $\tau_{EE/II}^{c}=-\tau_{EI/IE}^{c}$ instead leads to an increase of $\chi_{II}$ with the strength of chain motifs (brown line in Fig 11). We moreover predicted that setting $\tau_{EI/IE}^{c}=0$ and increasing $\tau_{EE/II}^{c}$ would lead to a decreasing $\chi_{II}$, and therefore a switch from a non-paradoxical to a paradoxical response (yellow line and frame in Fig 11). We directly verified this prediction in sparse networks by setting $\rho^{c}_{EI/IE}=c^{2}$ and therefore zero correlation coefficients $\tau_{EI/IE}^{c}=0$, while maintaining $\tau_{EE/II}^{c}=\tau^{c}>0$ and gradually increasing this value (Fig 11(b) and (c)).
We can alternatively use the effective connectivity to examine paradoxical responses in networks with heterogeneous variances and chain motifs. In this more general case, the explicit expression for $\mathbf{J}^{eff}$ is given by
$J_{ij}^{eff}=\left\{\begin{matrix}cJ+\alpha_{E}\sigma_{E}^{2}\tau_{EE}^{c}+\alpha_{I}\sigma_{I}\sigma_{E}\tau_{IE}^{c},&j\in E\\ -cgJ+\alpha_{E}\sigma_{E}\sigma_{I}\tau_{EI}^{c}+\alpha_{I}\sigma_{I}^{2}\tau_{II}^{c},&j\in I\end{matrix}\right.$ (56)
We find that this effective connectivity, predicts accurately the response functions in the full, sparse network (orange dashed lines in Figs. 910 and 11(b)).
We next investigated to which extent the stability of the E-E subnetwork of $\mathbf{J}^{eff}$ predicts paradoxical responses of the inhibitory population in the full network. The connectivity of the E-E subnetwork is given by
$\mathbf{J}_{EE}^{eff}=\left(cJ+\alpha_{E}\sigma_{E}^{2}\tau_{EE}^{c}+\alpha_{I}\sigma_{I}\sigma_{E}\tau_{IE}^{c}\right)\bm{e}\bm{e}^{\intercal},\,\,\,\bm{e}\in\mathbb{R}^{N_{E}}$ (57)
and has only one non-zero eigenvalue:
$\lambda_{EE}^{eff}=N_{E}(cJ+\alpha_{E}\sigma_{E}^{2}\tau_{EE}^{c}+\alpha_{I}\sigma_{I}\sigma_{E}\tau_{IE}^{c}).$ (58)
For the sparse network with a uniform occurrence $\rho^{c}$ we have $\tau^{c}=\tau_{EE/II}^{c}=-\tau_{EI/IE}^{c}$ (Figs. 9 and 10), so that the eigenvalue is
$\lambda_{EE}^{eff}=N_{E}cJ+\alpha_{E}^{2}N^{2}J^{2}c(1-c)\tau^{c}(1-g\gamma).$ (59)
With an initial condition of $N_{E}cJ<1$ and an inhibition-dominant network $(1-g\gamma<0)$, the eigenvalue remains negative as $\tau^{c}$ is increased, predicting correctly that the response of the inhibitory population remains non-paradoxical.
However, when setting $\tau_{EI/IE}^{c}=0$, the eigenvalue takes the form $\lambda_{EE}^{eff}=N_{E}cJ+\alpha_{E}^{2}N^{2}J^{2}c(1-c)\tau_{EE/II}^{c}$, and leads to a paradoxical inhibitory response as $\tau_{EE/II}^{c}$ is increased (Fig. 11(b), black triangle). Furthermore, this stability criterion for the E-E subnetwork accurately predicts the transition boundary between paradoxical and non-paradoxical responses in the $\tau_{EE/II}^{c}-\tau_{EI/EI}^{c}$ plane (Fig. 11(a)).
More generally, in the two-population networks we consider, chain motifs are specified by four independent parameters $\tau_{EE}^{c},\tau_{II}^{c},\tau_{EI}^{c}$ and $\tau_{IE}^{c}$. Figs. E.5-E.6 illustrate the effects on the low-rank approximation of varying each of these parameters independently. In particular, negative values of chain motifs can induce pairs of complex conjugate eigenvalues.
@@ V Discussion
Our findings demonstrate that chain motifs, a particular type of correlations between weights of pairs of synapses, exert a dominant impact on the dynamics in networks of recurrently connected neurons. Specifically, our mathematical analyses reveal that even a weak over-representation of chain motifs leads to the emergence of isolated eigenmodes distinct from networks with uncorrelated synaptic weights. In inhibition-dominated networks, these additional eigenmodes are associated with a positive eigenvalue that reflects positive feedback induced by chain motifs. This positive feedback leads to a potential instability that requires revisiting the classical balance conditions.
@@
The positive eigenmode generated by chain motifs strongly contributes to population-averaged responses to external inputs. In particular, we show that through this eigenmode, chain motifs can induce paradoxical responses in networks where such responses are not expected based on mean connectivity alone [10, 15, 19, 60, 61]. Conversely, in inhibition-stabilized networks that exhibit paradoxical responses in the absence of correlations between synapses, chain motifs can shift the dynamics towards non-paradoxical responses. Mean connectivity therefore does not predict the nature of the responses when chain motifs are present. Remarkably, we instead find that the response to external inputs is accurately predicted by a deterministic effective matrix $\mathbf{J}^{eff}$ (Eq. (50)), in which the mean synaptic strengths are corrected by the strength of chain and reciprocal motifs. In particular, paradoxical responses correspond to networks where the effective excitatory sub-network defined by $\mathbf{J}^{eff}$ is unstable by itself [15].
Our results for the impact of chain motifs on responses to inputs were therefore obtained by comparing two approaches: (i) a low-rank approximation based on the outliers in the spectrum of the full, random connectivity matrix $\mathbf{J}$; (ii) an approximation of the response using an effective deterministic connectivity $\mathbf{J}^{eff}$. The two approaches provide complementary pictures and have different strengths and weaknesses. The low-rank approximation can in principle be applied to networks with arbitrary connectivity structure. It preserves the dominant eigenvalues of the full connectivity, and is therefore expected to capture transient dynamics. However, it relies on the knowledge of the statistics of eigenvectors, which are in general not accessible analytically. The effective connectivity approach instead directly maps a large, random network with correlated synapses onto a classical deterministic network where each population is represented by one variable. A key difference with previous studies is that the connectivity weights in the effective deterministic network correspond to combinations of mean connectivity and motif strengths in the original random network. It is important to note that the effective connectivity matrix does not accurately capture the outliers in the eigenspectrum of the full connectivity matrix (Appendix D). While it predicts correctly the steady-state response, in general the effective connectivity is not expected to reproduce transient dynamics. Moreover, its form is specific to a connectivity structure based on pair-wise motifs.
An over-representation of chain motifs means that neurons with stronger inputs also have stronger outputs than expected by chance. This is precisely the case in the fully connected networks we investigated, where synaptic weights are continuous. In sparse networks, all synapses of a given type have the same strength, hence an over-representation of chain motifs means that neurons that receive more incoming synapses send out more outgoing synapses. This pattern is related, but not equivalent to network hubs. The concept of ‘hubs’ used in network science is generally based on the degree of individual nodes in a network [62]. In many examples, including biological neuronal networks, the degree distribution is heavy-tailed, meaning that a small number of nodes exhibit disproportionately high connectivity, and such nodes are called ‘hubs’. An over-representation of chain motifs implies a heavy-tailed degree distribution and therefore the existence of hub neurons with both high in-degree and high out-degree.
@@
Previous works have examined the effects of connectivity motifs on the continuous bulk of the spectrum [49, 32, 30, 31, 63]. Here we instead focus on eigenmodes corresponding to eigenvalues that lie outside of the bulk. While the abundance of all four motif types along with the variance were found to significantly impact the spectral radius of the bulk [30, 32], our results show that connectivity motifs have very distinct effects on the outliers, with chain motifs most dominantly shaping their behavior and determining the steady-state responses to external inputs.
Each connectivity eigenvalue is associated to a corresponding eigenvector and shapes the dynamics in this direction in the N-dimensional space of neurons. Since eigenvalues are densely packed in the connectivity bulk spectrum, many of these directions a priori play an important role in shaping the collective dynamics of neural circuits. However, when the spectral radius of the bulk approaches the line of instability at $Re(\lambda)=1$, the most positive eigenvalues close to instability dominate the dynamics, yielding a dimensionality of neural activity that is much lower than the number of neurons in the circuit [30]. The impact of a mode in the bulk to the overall network activity thereby continuously degrades with distance to the instability line, as reflected by smooth covariance spectra [32]. Outlier eigenvalues instead each determine a one-dimensional activity component that is weighing much differently for the overall network behavior due to the separation to the other eigenvalues. Their associated dynamics can thus be captured by a low-rank approximation and identified by a larger gap in the covariance spectrum.
@@
For the sake of analytical tractability, our study is based on a number of simplifying assumptions. We considered only linear, rate-based dynamics, while previous works outlined the effects of the non-linearity on paradoxical responses, in particular in supralinear-stabilized networks of excitatory and inhibitory neurons [14, 13, 12, 64, 65]. Although a full non-linear analysis of N-dimensional networks with correlated synapses is in general untractable, networks with low-rank connectivity allow for an exact low-dimensional reduction even in the presence of non-linearities [42, 66, 67]. Applying the low-rank approximation of correlated connectivity to supralinear stabilized networks is therefore an interesting direction for future research. Beyond rate-based models, numerical simulations of integrate-and-fire networks with low-rank connectivity structure suggest that the analytical insights from rate networks extend to spiking networks [68].
@@
How large is a given value of $\tau^{c}$? To get an intuitive understanding, it is useful to focus on sparse networks. In that case, the correlation coefficient $\tau^{c}$ is directly related to the occurrence probability $\rho^{c}$ (Eq. (40), which corresponds to the fraction of chain motifs observed across all possible chain motifs. A value of $\tau^{c}=0.1$ together with the connectivity probability $c=0.2$, corresponds to an occurrence probability of $\rho^{c}=0.056$. This is only $0.016$ above the chance level given by $c^{2}=0.04$. This implies that in an experiment where we examine $1000$ neuron triplets that could potentially form a chain motif, only about $56$ triplets will have the pair of synapses connected. Among these, $40$ triplets are expected purely due to chance, so only $16$ triplets contribute to an over-representation beyond chance. This level of over-representation therefore corresponds to a relatively small number of chain motifs, and yet has a substantial impact on the network dynamics.
Experimental studies of optogenetic perturbations have revealed the presence of paradoxical inhibitory responses throughout the cortex [19, 20, 21]. One proposed explanation is that these paradoxical responses stem from strong average recurrent excitation. Our findings show that, alternatively, paradoxical responses could be due to a weak over-representation of chain motifs. In recent physiological datasets, the measured strength of chain motifs is in the range of $\tau^{c}=0.05-0.1$ [30], and our results show that these values are compatible with paradoxical responses even if average recurrent excitation is weak. Additional investigations will however be needed to distinguish experimentally between these two possible mechanisms.